# ImFace++: A Sophisticated Nonlinear 3D Morphable Face Model with Implicit Neural Representations

### v2

@@ 4.2 Implementation Details
Network Architecture. All Mini-Net blocks ${\bm{\psi}}_{n}$ used by ExpNet $\mathcal{E}$, IDNet $\mathcal{I}$, and TempNet $\mathcal{T}$, are implemented as Multi-Layer Perceptrons (MLPs) with 3 hidden layers, each consisting of 32-dimensional hidden features. In contrast, the Mini-Net block utilized by DetailNet $\mathcal{D}$ is implemented as MLPs with 3 hidden layers, comprising 64-dimensional hidden features. For the activation function, we employ sine activation in every Mini-Net $\psi_{n}$, and the parameters are initialized as described in [45]. The frequency hyperparameter $w$ in the sine activation function determines the capability of the neural network. We use a smaller value of $w=30$ for the disentangled deformation fields and a larger value of $w=60$ for DetailNet to enhance the capability of the networks. At the end of each Mini-Nets block, the Fusion Network is implemented as a 128-dimensional, 3-layer MLP with a softmax activation function. During training, the parameters of the $k$ Mini-Nets are generated by $k$ corresponding Hyper Nets. This approach is adopted to achieve a more expressive latent space, following the insights from [23] and [64]. The Hyper Nets, denoted as $\phi_{n}$, consist of 1-layer MLPs activated by the Rectified Linear Unit (ReLU) function, with the hidden layer dimensionality set to 32. The Landmark-Nets $\eta$ and $\eta^{\prime}$ are composed of three 256-dimensional fully connected layers. The identity, expression, and detail embeddings are each represented as 160-dimensional vectors.
Training Settings. The trade-off parameters $\lambda_{1},\cdots,\lambda_{9}$ to train the networks are set to $33e$ 3, $21e$ 2, $15e$ 1, $51e$ 5, $31e$ 3, $31e$ 3, $21e$ 2, $21e$ 2, and $41e$ 4, respectively. The model is trained with Adam in an end-to-end manner. We train the model for 1,500 epochs with a learning rate of 0.0001. The training phase takes around 4 days on 4 32G NVIDIA V100 GPUs with minibatches of size 48. We randomly sample $m=4096$ points from the total 13,008 points in each batch to train the landmark consistency loss. During testing, it takes about 8 hours to optimize 319 samples on a single GPU. We use Marching Cubes[86] to reconstruct facial surfaces from the signed distance field, where the voxel resolution is set to $256^{3}$. All the meshes are rendered by Pyrender[96].
@@ 4.3.1 Reconstruction
We employ the proposed ImFace++ model to fit face scans by optimizing Eq. (21) and then compare the reconstruction results with several existing methods, including FLAME[6], FaceScape[41], the geometry model of i3DMM[33], NPHM[66], and the previous version, ImFace[87], which represent the current state-of-the-art. The official code of FLAME is used to fit the full face scans in the test set, utilizing 300 identity parameters and 100 expression parameters. For FaceScape, we use their released bilinear model built from 938 individuals for testing. The model parameters consist of 300 identity parameters and 52 expression parameters. It is important to note that our test scans are part of the training set for FaceScape. Furthermore, we modify their official code to fit full scans, rather than just landmarks, for improved results. For i3DMM and ImFace, we retrain these models on the same training set as used for ImFace++ to ensure a fair comparison. In both cases, the identity and expression embeddings are set to 160 dimensions, which aligns with our settings. For NPHM, we utilize the model trained with the official code. To maintain consistency with our study where we focus only on the facial region, 22 facial anchors are utilized and the parameter $K_{symm}$ in NPHM is set to 8.
@@
[…]
Figure 6: Reconstruction comparison with i3DMM[33], FLAME[6], FaceScape[41], NPHM[66], and ImFace[87]. Each column corresponds to a person with a non-neutral expression. Visually inspected, ImFace++ captures much richer shape variations and high-frequency details with more compact latent embeddings. Zoom in for a better view.
Figure 7: Comparison of correspondences with i3DMM[33], ImFace[87], and NPHM [66]. The leftmost face is morphed into multiple expressions (middle columns) and different identities (rightmost columns). ImFace++ surpasses the other backward-deformation methods, i.e. i3DMM and ImFace, and achieves results comparable to the forward-deformation method NPHM. Due to the forward-deformation mechanism, there are no cross-identity correspondence results for NPHM. Zoom in for a better view.
Qualitative Evaluation. Figure 6 provides a visual demonstration of the reconstruction results obtained by various models, with each row corresponding to a test individual displaying a non-neutral expression. i3DMM, being the first deep implicit model for human heads, exhibits limitations in capturing complex deformations and high-frequency details when faced with relatively intricate scenarios involving unseen facial shapes, resulting in artifacts on the reconstructed faces. FLAME and FaceScape, while effective in certain aspects, are constrained by fixed resolution and uniform mesh topology. FLAME excels at representing identity characteristics but struggles with non-linear deformations, leading to rigid facial expressions. FaceScape delivers more favourably results, primarily owing to high-quality training scans and the inclusion of test faces in the training set. However, it still struggles to accurately portray expression morphs and high-frequency details. NPHM utilizes local fields to model the SDF in a neutral expression, but the globally conditioned deformation field makes it challenging to represent fine-grained deformations and high-frequency details. Comparatively, ImFace excels in reconstructing facial shape by using the disentangled deformation fields, preserving subtle and rich nonlinear facial muscle deformations with fewer identity and expression latent parameters, such as frowns and pouts. However, it falls short of representing high-frequency details. In contrast, ImFace++ significantly improves upon its predecessor. The introduction of RDF within the template space enables ImFace++ to effectively handle intricate features, as demonstrated by its success in representing details such as eyelids (visible in the first two rows in Figure 6) and wrinkles (visible in the third to fifth rows).
@@
Quantitative Evaluation. To ensure a fair comparison, all faces in this evaluation undergo preprocessing to remove inner structures like eyeballs. This step allows for the computation of quantitative metrics within the same facial region for all models. Specifically, the symmetric Chamfer distance, F-score, and normal consistency (N.C.) are employed as metrics, with a strict F-score threshold set at 1mm. The results are presented in Table I, clearly demonstrating that ImFace++ outperforms its counterparts by a significant margin across all metrics, unequivocally validating its effectiveness. Additionally, a color-coded distance (fit-to-scan) is used to visualize the reconstruction errors in Figure 6. Overall, ImFace++ achieves more accurate face reconstruction results compared to other methods, particularly excelling in capturing nonlinear deformations and high-frequency facial details.
Metrics EDE ($mm$${}^{{\dagger}}$ TDE ($mm$${}^{{\dagger}}$
@@
TABLE II: Quantitative correspendences comparison with the state-of-the-art methods. (${}^{{\dagger}}$ower is better.)
4.3.2 Correspondence
In contrast to existing methods that typically rely on accurate face registration, ImFace++ leverages the disentangled deformation fields to automatically learn correspondences among different face scans. Furthermore, we incorporate the training critic, specifically the landmark consistency loss ${\cal L}_{lmk_{c}}^{i}$, to enhance this feature. This evaluation is designed to assess the correspondence establishment capabilities of various INR-based models. In particular, given two 3D faces, we use i3DMM, ImFace, and ImFace++ to fit them, subsequently deforming the densely sampled points to the template space. This process ensures point-to-point correspondences, which are achieved through a nearest neighbor search using the KD tree algorithm for efficiency. Considering that NPHM is a forward-deformation-based method, we deform the sampled points in a neutral expression to another expression and examine the correspondences.
Qualitative Evaluation. Figure 7 provides a visual comparison of the results, with manually painted color patterns on the shapes to facilitate quality assessment. In alignment with our approach, i3DMM incorporates a landmark consistency loss to enhance correspondences. However, it employs a single deformation field to represent both expression and identity deformations in an intertwined manner. Consequently, the correspondences may not be accurately learned in certain areas, such as the mouth (regions F4-F5 in the painted color patterns), forehead (regions B3-B6), and nostril (regions E4-E5). ImFace successfully establishes correspondences across various expressions and identities. Nonetheless, it is noticeable that minor internal texture discrepancies occasionally occur around the mouth (regions F4-F5, G4-G5). This is primarily because facial shapes undergo significant changes in these localized areas with different expressions. ImFace++ addresses this challenge by employing densely sampled landmarks (increasing from 68 in ImFace to 4,096 in ImFace++) to supervise landmark consistency loss during training, resulting in more accurate outcomes. Due to the forward-deformation pipeline, NPHM naturally exhibits more consistent correspondences than backward-deformation approaches. However, it may overlook exaggerated expressions like an open mouth, potentially leading to semantically incorrect correspondences (regions G4-G5). In contrast, ImFace++ achieves comparable or even superior results to NPHM, demonstrating its effectiveness.
@@ 4.3.2 Correspondence
Figure 8: Module analysis. The disentangled deformation fields are responsible for learning deformations related to facial expressions and identities. The refinement displacement field enables the representation of high-frequency details.
@@
Quantitative Evaluation. To further validate the effectiveness of our method, we conduct a quantitative evaluation on the learned correspondences similar to i3DMM [33]. For each test face $F$ with identity $i$ from FaceScape, we first identify the face $F_{ne}$ with the same identity but in a neutral expression. We then upsample the registered mesh resolution and sample 22,481 points from faces $F$, $F_{ne}$, and the template face, which are not used during training, serving as the ground-truth annotations for correspondences. We use i3DMM, ImFace, and ImFace++ to deform the points sampled from face $F$ to the canonical space and the template space, respectively. We calculate the average Expression Deformation Error (EDE), which quantifies the errors between the deformed points and the ground-truth points in the canonical space, and the Total Deformation Error (TDE), which measures the errors in the template space. Note that i3DMM employs a single deformation field to represent both expression and identity deformations, allowing it to compute only the TDE metric. On the other hand, NPHM models only the deformation of expression and we can calculate only the EDE metric. For NPHM, we follow the approach of SNARF [31] and utilize iterative root finding to locate the canonical positions of the points sampled from face $F$, allowing us to compute the EDE metric. As shown in Table II, ImFace++ outperforms the other methods in terms of correspondence quality. Notably, it even surpasses the forward-deformation method, NPHM, underscoring the effectiveness of the disentangled deformation fields, which play a crucial role in addressing the challenges associated with learning the RDF. It is worth mentioning that the results in Table II also encompass the reconstruction errors.
@@ 4.4 Representation Ability
To further assess the representation ability of ImFace++, we conduct comprehensive evaluations in both the observed and latent representation spaces. These evaluations comprise three main aspects. i) Module Analysis: we analyze the learned content of the various modules introduced in ImFace++. ii) Compact Embedding Space Interpolation: given the compact embedding space of ImFace++, we perform animations on the generated or reconstructed facial shapes through interpolation in the identity, expression, and detail embeddings. iii) Variation Visualization: we apply Principal Components Analysis (PCA) to the learned expression and identity embeddings to visualize model variations. Additionally, we employ t-SNE [85] to assess the unsupervised learning ability of ImFace++.
@@
Module Analysis. Figure 8 presents the reconstructed test scans by ImFace++, offering insights into the content learned by the different modules proposed in our method. In the third and fourth columns, we can observe that the disentangled deformation fields, managed by ExpNet for expression and IDNet for identity, successfully capture rich and complex deformations. However, the reconstructed facial shape in these cases lacks high-frequency details. Nevertheless, after training the RDF within the template space, ImFace++ is capable of representing facial shapes that incorporate personalized high-frequency details. This is particularly evident in features like the eyelids in the first two rows and the mouth shape in the third row. In the ablation study section, we offer further evidence and a more comprehensive analysis.
@@
Compact Embedding Space Interpolation. The seamless animation of facial shapes through interpolation in the identity, expression, and detail embeddings demonstrates that ImFace++ has learned compact latent embedding spaces. As illustrated in Figure 9, our identity embeddings encompass the overall shape, including the positioning of the nose, eyes, and mouth, as well as the gender of the face. In contrast, the expression embeddings capture deformations specific to facial components. Furthermore, when comparing the first and last columns in Figure 9, it becomes evident that the detail embeddings represent personalized high-frequency details of the face, such as wrinkles and dimples. As demonstrated, even when the embeddings used for interpolation exhibit significant changes over a wider range (in the last two rows), our method is capable of generating smooth and coherent results.
Figure 9: Embedding interpolation results. The Trainset and Testset scans can be smoothly animated by interpolating in the embedding space.
Figure 10: Visualization of the expression and identity principal components of ImFace++ on FaceScape [41].
Figure 11: Visualizing the distributions of high-dimensional expression and identity embeddings using t-SNE. (a) The embedding distribution of 6 typical expressions from the training set. (b) The identity embedding distribution from the test set, which involves 16 individuals.
@@
Variation Visualization. We apply PCA to the learned expression and identity embeddings to visualize the model capacity, as depicted in Figure 10. The standard deviations for the expression and identity embeddings are set to ±$\pm~{}0.25$ and ±$\pm~{}5.0$, respectively. The embeddings encoding facial details are fixed at the mean value calculated from the entire training set. In particular, four expression principal components are visualized in Figure 10 (a). Despite significant expression changes, the faces maintain a consistent identity. ImFace++ is capable of producing vivid expressions by learning expression components from thousands of unique embeddings. In Figure 10 (b), a similar phenomenon is observed in the learned identity components, where facial expressions remain stable while identity varies. These experimental results demonstrate that a good disentanglement between expression and identity has been achieved, which is crucial for generating novel faces by reweighting the singular values.
Additionally, we employ t-SNE[85] to visualize the learned high-dimensional expression embeddings from 4,992 training scans, encompassing six typical expressions, as depicted in Figure 11 (a). It is evident that our network can effectively distinguish different expression types through unsupervised learning, solely from expression-related shape morphs, without relying on explicit expression labels. Furthermore, Figure 11 (b) presents the visualization of identity embeddings, involving 319 face scans from 16 individuals in the test set. Upon visual inspection, our model successfully captures inter-subject differences even under various complicated expressions.
4.5 Ablation Study
ImFace++ is constructed upon the following core components: Disentangled Deformation Fields (DDF), Refinement Displacement Field (RDF), Neural Blend-Field (Blend), and Improved Expression Embedding Learning (Extend). We experimentally verify the credits of these designs. When performing ablation of RDF, the model is essentially downgraded to ImFace. During its training process, we increase the number of the sampled points to 4,096 to supervise the landmark consistency loss, ensuring a fair comparison with ImFace++.
@@ 4.5 Ablation Study
TABLE III: Quantitative ablation study results. $*{}^{*}$ denotes the use of 4,096 landmarks to supervise the landmark consistency loss during model training.

### v3

@@ 4.2 Implementation Details
Network Architecture. All Mini-Net blocks used by ExpNet, IDNet, and TempNet, are implemented as Multi-Layer Perceptrons (MLPs) with 3 hidden layers, each consisting of 128-dimensional hidden features, while the Mini-Net block utilized by DetailNet is implemented as MLPs with 3 hidden layers, comprising 256-dimensional hidden features. For the activation function, we employ sine activation in every Mini-Net, and the parameters are initialized as described in [45]. The frequency hyperparameter $w$ in the sine activation function determines the capability of the neural network. We use a smaller value of $w=30$ for the disentangled deformation fields and a larger value of $w=60$ for DetailNet to enhance the capability of the networks. At the end of each Mini-Nets block, the Fusion Network is implemented as a 128-dimensional, 3-layer MLP with a softmax activation function. During training, the parameters of Mini-Nets are generated by corresponding Hyper Nets. This is adopted to achieve a more expressive latent space, following the insights from [23] and [64]. The Hyper Nets consist of 1-layer MLPs activated by the Rectified Linear Unit (ReLU) function, with the hidden layer dimensionality set to 32. The Landmark-Nets are composed of three 256-dimensional fully connected layers. The identity, expression, and detail embeddings are individually represented as 160-dimensional vectors.
Training and Inference Settings. The trade-off parameters $\lambda_{1},\cdots,\lambda_{9}$ to train the networks are set to $33e$ 3, $21e$ 2, $15e$ 1, $51e$ 5, $31e$ 3, $31e$ 3, $21e$ 2, $21e$ 2, and $41e$ 4, respectively. The model is optimized with Adam in an end-to-end manner. We train the model for 1,500 epochs with a learning rate of 0.0001. The training phase takes around 4 days on 4 32G NVIDIA V100 GPUs with minibatches of size 48. We randomly sample $m=4096$ points from the total 13,008 points in each batch to train the landmark consistency loss. During testing, it takes about 8 hours to infer 319 samples on a single GPU. We use Marching Cubes [86] to reconstruct facial surfaces from the signed distance field, where the voxel resolution is set to $256^{3}$. All the meshes are rendered by Pyrender[110].
Figure. 6: Reconstruction comparison with i3DMM[33], FLAME[6], FaceScape[41], NPHM[66], and ImFace[87]. Each column corresponds to a person with a non-neutral expression. Visually inspected, ImFace++ captures much richer shape variations and high-frequency details with more compact latent embeddings. Zoom in for a better view.
@@ 4.3.1 Reconstruction
We employ the ImFace++ model to fit face scans by optimizing Eq. (21) and then compare the reconstruction results with several existing methods, including FLAME [6], FaceScape [41], the geometry model of i3DMM [33], NPHM [66], and our previous version, ImFace [87], which represent the current state-of-the-art. The official code of FLAME is used to fit the full face scans in the test set, utilizing 300 identity parameters and 100 expression parameters. For FaceScape, we use their released bilinear model built from 938 individuals for testing. The model parameters consist of 300 identity parameters and 52 expression parameters. It is important to note that our test scans are part of the training set for FaceScape. Furthermore, we modify their official code to fit full scans, rather than just landmarks, for improved results. For i3DMM and ImFace, we retrain these models on the same training set as used for ImFace++ to ensure fair comparison. In both cases, the identity and expression embeddings are set to 160 dimensions, which aligns with our settings. For NPHM, we utilize the model trained with the official code. To maintain the consistency with our study where we focus only on the facial region, 22 facial anchors are utilized and the parameter $K_{symm}$ in NPHM is set to 8.
Qualitative Evaluation. Fig. 6 provides a visual demonstration of the reconstruction results obtained by various models, with each row corresponding to a test individual displaying a non-neutral expression. i3DMM, the first deep implicit model for human heads, exhibits limitations in capturing complex deformations and high-frequency details when confronting with relatively intricate scenarios involving unseen facial shapes, incurring artifacts on the reconstructed results. FLAME and FaceScape, while effective in certain aspects, are constrained by fixed resolutions and uniform mesh topologies. FLAME excels at representing identity characteristics but struggles with non-linear deformations, leading to rigid facial expressions. FaceScape delivers more favourably results, primarily owing to high-quality training scans and the inclusion of test faces in the training set. However, it still fails to accurately portray expression morphs and high-frequency details. NPHM utilizes local fields to model the SDF in a neutral expression, but the globally conditioned deformation field makes it challenging to represent fine-grained deformations and high-frequency details. Comparatively, ImFace excels in reconstructing facial shape by using the disentangled deformation fields, preserving subtle and rich nonlinear facial muscle deformations with fewer identity and expression latent parameters, such as frowns and pouts. However, it falls short of representing high-frequency details. In contrast, ImFace++ significantly improves upon its predecessor. The introduction of RDF within the template space enables it to effectively handle intricate features, as demonstrated by its success in representing details such as eyelids (visible in the first two rows in Fig. 6) and wrinkles (visible in the third to fifth rows).
Additionally, we select 100 facial shapes from the LYHM dataset for cross-dateset evaluation. We calculate the rotation, translation, and scale for each shape using 3D landmarks. We then register these facial shapes to our coordinate system and preprocess them as described in Sec. 4.1. The evaluation is performed using the model trained on the FaceScape dataset. We present the cross-dataset results in Fig. 7, demonstrating that our model generalizes well to new data.
Figure. 7: Cross-dataset reconstruction results. The model is trained on FaceScape and tested on LYHM. The Chamfer distance for each method is listed below.
@@
TABLE I: Quantitative reconstruction comparison with the state-of-the-art methods (Lower is better; Higher is better).
@@
Quantitative Evaluation. To ensure fair comparison, all faces in this evaluation undergo preprocessing to remove inner structures like eyeballs. This step allows for the computation of quantitative metrics within the same facial region for all models. Specifically, the symmetric Chamfer distance, F-score, and normal consistency (N.C.) are employed as metrics, with a strict F-score threshold set at 1mm. The results are presented in Tab. I, clearly demonstrating that ImFace++ outperforms its counterparts by a significant margin across all the metrics, unequivocally validating its effectiveness. Additionally, a color-coded distance (fit-to-scan) is used to visualize the reconstruction errors in Fig. 6. Overall, ImFace++ achieves more accurate face reconstruction results compared to other methods, particularly excelling in capturing nonlinear deformations and high-frequency facial details.
4.3.2 Correspondence
In contrast to existing methods that typically rely on accurate face registration, ImFace++ leverages the disentangled deformation fields to automatically learn correspondences among different face scans. Furthermore, we incorporate the training critic, i.e. the landmark consistency loss ${\cal L}_{lmk_{c}}^{i}$, to enhance the feature. This evaluation is designed to assess the correspondence establishment capabilities of various INR-based models. In particular, given two 3D faces, we use i3DMM, ImFace, and ImFace++ to fit them, subsequently deforming the densely sampled points to the template space. This process ensures point-to-point correspondences, which are achieved through a nearest neighbor search using the KD-Tree algorithm for efficiency. Considering that NPHM is a forward-deformation-based method, we deform the sampled points in a neutral expression to another expression and examine the correspondences.
Qualitative Evaluation. Fig. 8 provides visual comparison of the results, with manually painted color patterns on the shapes to facilitate quality assessment. This is achieved by mapping the colors from a source face to the target one utilizing the established point-to-point correspondence. In alignment with our approach, i3DMM incorporates a landmark consistency loss to enhance correspondences. However, it employs a single deformation field to represent both expression and identity deformations in an intertwined manner. Consequently, the correspondences may not be accurately learned in certain areas, such as the mouth (regions F4-F5 in the painted color patterns), forehead (regions B3-B6), and nostril (regions E4-E5). ImFace successfully establishes correspondences across various expressions and identities. Nonetheless, it is noticeable that minor internal texture discrepancies occasionally occur around the mouth (regions F4-F5, G4-G5). This is primarily because facial shapes undergo significant changes in these localized areas with different expressions. ImFace++ addresses this challenge by employing densely sampled landmarks (increasing from 68 in ImFace to 4,096 in ImFace++) to supervise the landmark consistency loss during training, resulting in more accurate outcomes. Due to the forward-deformation pipeline, NPHM naturally exhibits more consistent correspondences than backward-deformation approaches. However, it may overlook exaggerated expressions like a closed eye, potentially leading to semantically incorrect correspondences (regions G4-G5). In contrast, ImFace++ achieves comparable or even superior results to NPHM, demonstrating its effectiveness.
Metrics EDE ($mm$) TDE ($mm$)
@@ 4.3.2 Correspondence
TABLE II: Quantitative correspendence comparison with the state-of-the-art methods. (Lower is better.)
Figure. 8: Comparison of correspondences with i3DMM[33], ImFace[87], and NPHM [66]. The leftmost face is morphed into multiple expressions (middle columns) and different identities (rightmost columns). ImFace++ surpasses the other backward-deformation methods, i.e. i3DMM and ImFace, and achieves the results comparable to the forward-deformation method NPHM. Due to the forward-deformation mechanism, there are no cross-identity correspondence results for NPHM. Zoom in for a better view.
@@
Figure. 9: Module analysis. The disentangled deformation fields are responsible for learning deformations related to facial expressions and identities. The refinement displacement field enables the representation of high-frequency details.
@@
Quantitative Evaluation. To further validate the effectiveness of our method, we conduct quantitative evaluation on the learned correspondences similar to i3DMM [33]. For each test face $F$ with identity $i$ from FaceScape, we first identify the face $F_{ne}$ with the same identity but in a neutral expression. We then upsample the registered mesh resolution and sample 22,481 points from faces $F$, $F_{ne}$, and the template face, which are not used during training, serving as the ground-truth annotations for correspondences. We use i3DMM, ImFace, and ImFace++ to deform the points sampled from $F$ to the identity space and the template space, respectively. We calculate the average Expression Deformation Error (EDE), which quantifies the errors between the deformed points and the ground-truth points in the identity space, and the Total Deformation Error (TDE), which measures the errors in the template space. Note that i3DMM employs a single deformation field to represent both expression and identity deformations, allowing it to compute only the TDE metric. On the other hand, NPHM models only the deformation of expression and we can calculate only the EDE metric. For NPHM, we follow the approach of SNARF [31] and utilize iterative root finding to locate the canonical positions of the points sampled from $F$, allowing us to compute the EDE metric. As shown in Tab. II, ImFace++ outperforms the other methods in terms of correspondence quality. Notably, it even surpasses the forward-deformation method, NPHM, underscoring the effectiveness of the disentangled deformation fields, which play a crucial role in addressing the challenges associated with learning the RDF. It is worth mentioning that the results in Tab. II also encompass the reconstruction errors.
Figure. 10: Embedding interpolation results. The Trainset and Testset scans can be smoothly animated by interpolating in the embedding space.
Figure. 11: Visualization of the expression and identity principal components of ImFace++ on FaceScape [41].
Figure. 12: Visualizing the distributions of high-dimensional expression and identity embeddings using $t$-SNE. (a) Embedding distribution of 20 typical expressions from the training and testing sets. (b) The identity embedding distribution from the test set, which involves 16 individuals.
@@ 4.4 Representation Ability
To further assess the representation ability of ImFace++, we conduct comprehensive evaluation in both the observed and latent representation spaces. The evaluation comprises three main aspects. i) Module Analysis: we analyze the learned content of the various modules introduced in ImFace++. ii) Compact Embedding Space Interpolation: given the compact embedding space of ImFace++, we perform animations on the generated or reconstructed facial shapes through interpolation in the identity, expression, and detail embeddings. iii) Variation Visualization: we apply Principal Components Analysis (PCA) to the learned expression and identity embeddings to visualize model variations. Additionally, we employ $t$-SNE [85] to assess the unsupervised learning ability of ImFace++.
@@
Module Analysis. Fig. 9 presents the reconstructed test scans by ImFace++, offering insights into the content learned by the different modules proposed in our method. In the third and fourth columns, we can observe that the disentangled deformation fields, managed by ExpNet for expression and IDNet for identity, successfully capture rich and complex deformations. However, the reconstructed facial shapes in these cases lack high-frequency details. Nevertheless, after training the RDF within the template space, ImFace++ is capable of representing facial shapes that incorporate personalized high-frequency details. This is particularly evident in features like the eyelids in the first two rows and the mouth shape in the third row. In the ablation study section, we offer further evidences and a more comprehensive analysis.
@@
Compact Embedding Space Interpolation. The seamless animation of facial shapes through interpolation in the identity, expression, and detail embeddings demonstrates that ImFace++ is able to learn compact latent embedding spaces. As illustrated in Fig. 10, our identity embeddings encompass the overall shape, including the positioning of the nose, eyes, and mouth, as well as the gender of the face. In contrast, the expression embeddings capture deformations specific to facial components. Furthermore, when comparing the first and last columns in Fig. 10, it becomes evident that the detail embeddings represent personalized high-frequency details of the face, such as wrinkles and dimples. As demonstrated, even when the embeddings used for interpolation exhibit significant changes over a wider range (in the last two rows), our method is capable of generating smooth and coherent results.
@@
Variation Visualization. We apply PCA to the learned expression and identity embeddings to visualize the model capacity, as depicted in Fig. 11. The standard deviations for the expression and identity embeddings are set to ±$\pm~{}0.25$ and ±$\pm~{}5.0$, respectively. The embeddings encoding facial details are fixed at the mean value calculated from the entire training set. In particular, four expression principal components are visualized in Fig. 11 (a). Despite significant expression changes, the faces maintain a consistent identity. ImFace++ is capable of producing vivid expressions by learning expression components from thousands of unique embeddings. In Fig. 11 (b), a similar phenomenon is observed in the learned identity components, where facial expressions remain stable while identity varies. These experimental results demonstrate that a good disentanglement between expression and identity is achieved, which is crucial for generating novel faces by reweighting the singular values.
Additionally, we employ $t$-SNE[85] to visualize the learned high-dimensional expression embeddings from the whole training and testing scans, encompassing 20 expressions, as depicted in Fig. 12 (a). Some entanglement remains for certain expressions, such as sadness and anger, and this is primarily due to the inherent complexity of distinguishing these similar expressions. From the perspective of facial expression recognition, these expressions often share overlapping facial features, such as furrowed brows and downturned lips, making them difficult to differentiate. Despite this, our method shows promising capabilities in differentiating between various expression types through unsupervised learning, solely from expression-related shape morphs, without the need of explicit expression labels.
@@
TABLE III: Quantitative ablation study results. ImFace denotes using 4,096 landmarks to supervise the landmark consistency loss during model training.
Furthermore, Fig. 12 (b) presents the visualization of identity embeddings, involving 319 face scans from 16 individuals in the test set. Upon visual inspection, our model successfully captures inter-subject differences even under various complicated expressions.
4.5 Ablation Study
ImFace++ is constructed upon the following core components: Disentangled Deformation Fields (DDF), Refinement Displacement Field (RDF), Neural Blend-Field (Blend), Improved Expression Embedding Learning (Extend), and Hyper Net. We experimentally validate the effectiveness of these components, including evaluating the landmark consistency loss. When performing ablation of RDF, the model is essentially downgraded to ImFace. During its training process, we increase the number of the sampled points to 4,096 to supervise the landmark consistency loss, ensuring fair comparison with ImFace++.